OpenStudy (anonymous):

There is a 6% chance that a vaccine will cause a certain side effect. A number of patients are given the vaccine. We are interested in the number of patients vaccinated until the first side effect is observed. a) Define the random variable X. b) Verify that this setting is a geometric setting. c) Find the probability that the 5th patient is the first to experience a side effect. Construct a probability distribution table for X up through X=5. Please check with your instructor on how to submit your work d) How many patients would you expect to vaccinate before the first side effect is o

OpenStudy (anonymous):

observed?

OpenStudy (anonymous):

@saifoo.khan @Agent_Sniffles @Hero @phi @JamesJ anyone?

OpenStudy (jamesj):

So, which parts are you stuck on?

OpenStudy (anonymous):

the entire question lol

OpenStudy (anonymous):

I got (a). X = number of patients until the first side effect is observed

OpenStudy (jamesj):

Yes So if there is just one patient, what is the value of X?

OpenStudy (anonymous):

Ummm.... 1?

OpenStudy (jamesj):

btw, I assume this should read: "Construct a probability distribution table for X up through n [not X] = 5."

OpenStudy (anonymous):

how do I verify (b)

OpenStudy (jamesj):

So if there is one patient, then you expect one patient should suffer the side effect? No. What is the probability that that one patient suffers the side effect?

OpenStudy (anonymous):

Ummmmmm

OpenStudy (anonymous):

im not sure how to find that

OpenStudy (jamesj):

"There is a 6% chance that a vaccine will cause a certain side effect." Hence, if n = 1, X = ... ?

OpenStudy (anonymous):

6%

OpenStudy (anonymous):

?

OpenStudy (jamesj):

X(n=1) = 0.06, yes Now, if there are two patients, what is the probability that the second patient is the first one to experience the side effects? Pr(2nd patient has side effect) = Pr(2nd patient has side effect | 1st doesn't) Pr(1st does) Now, what are those two terms on the right side equal to?

OpenStudy (jamesj):

**Correction: Pr(2nd has side effect) = Pr(2nd has side effect | 1st doesn't) . Pr(1st doesn't)

OpenStudy (anonymous):

uhh

OpenStudy (jamesj):

Pr(2nd has side effect) = probability that 2nd patient is the first to suffer the side effect. Now, what is that equal to?

OpenStudy (jamesj):

If the 2nd patient is the first to suffer, then the 1st didn't. Hence Pr(2nd does) = Pr(2nd does given that the 1st doesn't) TIMES Pr(1st doesn't)

OpenStudy (jamesj):

i.e., Pr(2nd has side effect) = Pr(2nd has side effect | 1st doesn't) . Pr(1st doesn't) Make sense?

OpenStudy (jamesj):

Where did you go? Can't help you if don't engage in conversation.

OpenStudy (anonymous):

@JamesJ im back, sorry

OpenStudy (jamesj):

Right. Does this formula now make sense? Pr(2nd has side effect) = Pr(2nd has side effect | 1st doesn't) . Pr(1st doesn't)

OpenStudy (jamesj):

or being more explicit, Probability(2nd being first patient to have side effect) = Probability(2nd has side effect | 1st doesn't) . Probability(1st doesn't)

OpenStudy (anonymous):

Ok

OpenStudy (jamesj):

So what are the values of Probability(2nd has side effect | 1st doesn't) and Probability(1st doesn't suffer side effects)

OpenStudy (anonymous):

I have no clue how to find that

OpenStudy (anonymous):

if x=1, n=.06

OpenStudy (jamesj):

Well, Probability(1st doesn't suffer side effects) + Probability(1st does suffer side effects) = 1 What is Probability(1st does suffer side effects) ? And therefore what is Probability(1st doesn't suffer side effects) ?

OpenStudy (anonymous):

Its P(0.06?

OpenStudy (jamesj):

Probability(1st does suffer side effects) = 0.06, yes So what is Probability(1st doesn't suffer side effects) = ... ?

OpenStudy (anonymous):

.96?

OpenStudy (jamesj):

No Probability(1st doesn't suffer side effects) = 1 - 0.06 = 0.94 Make sense?

OpenStudy (anonymous):

oh yeah thats what i meant oops lol

OpenStudy (jamesj):

Now, we want to calculate Probability(2nd being first patient to have side effect) = Probability(2nd is first to have side effect | 1st doesn't) . Probability(1st isn't first) Now, think! What is Probability(2nd is first to have side effect | 1st doesn't) ?

OpenStudy (jamesj):

Probability(2nd is first to have side effect | 1st doesn't) is just the same as Probability(2nd having side effect) as the events are what's called independent events. What is the value of Probability(2nd having side effect) ?

OpenStudy (anonymous):

ugh im so confused

OpenStudy (jamesj):

Just stay with me. What is confusing you right now?

OpenStudy (anonymous):

Probability(2nd is first to have side effect | 1st doesn't)

OpenStudy (anonymous):

can u just explain to me how to find that

OpenStudy (jamesj):

This means what is the probability that the 2nd patient is the first to have the side effect given that (that is what "|" means, "|" = "given that") the first patient hasn't suffered the side effects. Now Probability(2nd is first to have side effect | 1st doesn't) is just equal to the Probability(2nd patient has side effects) because whether or not the 1st patient suffers side effects has nothing to do with whether or not the second patient suffers side effects. The drug effects each patient independently, just like separate rolls of dice. Yes? If so, what is Probability(2nd patient has side effects) = .... ?

OpenStudy (jamesj):

Well, the probability that ANY patient suffers side effects is 6% or 0.06 Hence Probability(2nd patient has side effects) = 0.06

OpenStudy (jamesj):

Therefore Probability(2nd being first patient to have side effect) = Probability(2nd has side effect | 1st doesn't) x Probability(1st doesn't) = 0.06 x 0.94 Compressing notation now write Probability(2nd being first patient to have side effect) as Pr(2 first) = 0.06 x 0.94 Yes?

OpenStudy (anonymous):

Ok!

OpenStudy (jamesj):

So what is Pr(3 first) = ... ? Expand it out using the | notation

OpenStudy (anonymous):

.06 x .94 x ....

OpenStudy (jamesj):

Write it out as ... Pr(3 first) = Pr(3 first | not 1 first and not 2 first) . Pr(not 1 first) . Pr(not 2 first) = what . what . what

OpenStudy (anonymous):

confused, sorry

OpenStudy (jamesj):

The probability that the 3rd patient is the first to suffer side effects is equal to (The probability that the 3rd patient suffers sides effects GIVEN that the 1st and 2nd patients haven't suffered side effects) TIMES (The probability that the first patient didn't suffer side effects) TIMES (The probability that the second patient didn't suffer side effects) In notation, that is just Pr(3 first) = Pr(3 first | not 1 first and not 2 first) . Pr(not 1 first) . Pr(not 2 first)

OpenStudy (anonymous):

Im gonna do this prob later i gtg bye